OpenStudy (anonymous):

Solve. 64^x = 16^x–1

OpenStudy (lgbasallote):

64 is 4^3..16 is 4^2 so... (4^3)^x = (4^2)^(x-1) 4^3x = 4^(2)(x-1) i'll leave the rest to you :DDD

OpenStudy (anonymous):

64^x = 16^(x - 1) (2^6)^x = (2^4)^(x - 1) Exponent rules. 2^(6x) = 2^(4x - 4) Log base 2 of both sides. 6x = 4x - 4 2x = -4 x = -2